OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x - 9

OpenStudy (anonymous):

5x-4x=-9-2

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

NO :(

OpenStudy (anonymous):

why? just add them

OpenStudy (anonymous):

5x-4x=x and -9-2=-11

OpenStudy (anonymous):

so,x=-11

OpenStudy (anonymous):

x = -11 5x + 2 = 4x − 9 Subtract 4x from both Sides -----> x + 2 = -9 Subtract 2 from both sides x = -11